# 2cm2cm1.5cm1.5cm

### v1

@@ Abstract.
Finding a largest Euclidean ball in a given convex body $K$ in $\mathbb{R}^{d}$, and finding the largest volume ellipsoid in $K$ are two problems of very different nature. The first is a Euclidean problem, where similar copies of the origin centered closed unit ball are considered, whereas in the second, we search among affine copies of $\mathbf{B}^{d}$. In this paper, we interpolate between the two problems by considering ellipsoids of revolution. We generalize the problem by looking at pairs of convex bodies $K$ and $L$ and the largest volume affine image of $K$ contained in $L$, where we make certain restrictions on the affine transformation.
@@ 1.2. A more general problem
A natural generalization of this question has been studied by Lewis [Lew79], V. Milman [TJ89, Theorem 14.5], Giannopoulos, Perissinaki, and Tsolomitis [GPT01], Bastero and Romance [BR02], and Gordon, Litvak, Meyer, and Pajor [GLMP04]: given convex bodies $K$ and $L$ in $\mathbb{R}^{d}$, find a maximal volume affine image of $K$ that is contained in $L$. In this work, we study the same problem under additional restrictions on the affine transformation.
@@
Note a subtle but very important difference between this general case and ellipsoids. To define ellipsoids of revolution with an $s$-dimensional axis, we only need to specify the dimension $s$ of the axis, but not the axis $F$. In the general case introduced here, we fix $F$, apply an $F$-operator $A$ and then allow a rotation $R$. The advantage of this is that in this way, we can specify what ‘part’ of $K$ we allow to be changed by an affine transformation (the $F$-component), and what part (the ${F}^{\bot}$ component) is rigid, that is, it only allows similarities. A simple illustration is the task of finding the largest area isosceles triangle contained in a planar set $L$. In our framework, we simply take $K$ to be a regular triangle, and $F$ to be one of its latitude lines. Then $\mathcal{R}_{F}\!\left(K\right)$ is exactly the set of isosceles triangles.
When $F=\{\mathbf{0}\}$, then Problem 4 is the problem discussed at the beginning of this subsection considered by Lewis and others. On the other hand, when $F=\mathbb{R}^{d}$, then we are looking for the largest similar copy of $K$ in $L$.
@@ 1.3. Results
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ with $K\subseteq L$. Following [GLMP04], we call $(v,p)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ a contact pair of $K$ and $L$ if $v\in\operatorname{bd}K\cap\operatorname{bd}L,$ and $p\in\operatorname{bd}{K}^{\circ}\cap\operatorname{bd}{L}^{\circ}$ normalized by the condition $\left\langle p,v\right\rangle=1$. In other words, $p$ is an outer normal vector of a common support hyperplane of $K$ and $L$ at $v$, with the appropriate normalization. We use $P_{F}$ to denote the orthogonal projection onto $F$.
@@
Let $K,L\subset\mathbb{R}^{d}$ be convex bodies such that $K\subseteq L$, and let $F$ be a linear subspace in $\mathbb{R}^{d}$. Assume that $K$ is a solution to John’s general problem with fixed axis $F$ for $K$ and $L$, that is, Problem 3. Then there is
@@
    a translation vector $z\in\mathrm{int}K$,
  2. (2)
    contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L-z$ and $K-z$,
@@
(1.1) $P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)P_{F}=P_{F};$
@@
(1.3) $\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}=\mathbf{0};$
(1.4) $\operatorname{tr}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)=\sum\limits_{i\in[m]}\upalpha_{i}=d.$
As a corollary, we obtain the following.
Corollary 1.5 (Containment for John’s general problem with fixed axis).
With the assumptions of Theorem 1.4 and the translation vector $z\in\mathrm{int}K$ found therein, we have
$(K-z)\cap F\subset\left(L-z\right)\cap F\subset-d\cdot P_{F}(K-z).$
In the case $F=\mathbb{R}^{d}$, Theorem 1.4 and Corollary 1.5 were obtained in [GLMP04].
@@
With the notation of Theorem 1.4, assume that $K$ is a solution to John’s general problem with rotated axis for $K$, $L$ and the subspace $F$. Then there is a translation vector $z\in\mathrm{int}K$, and there are contact pairs of $L-z$ and $K-z$, and weights as in Theorem 1.4 such that in addition to identities (1.1) – (1.4), we also have
(1.5) $\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}.$
@@
Let an origin-centered ellipsoid $E$ be a solution to Problem 1 for a convex body $K$ and a linear subspace $F$ in $\mathbb{R}^{d}$ of dimension $s\in\{0,1,\dots,d\}$. Then
@@
    $\operatorname{inrad}\left(K\cap{F}^{\bot}\right)\leq\frac{d}{d-s}\cdot\mathrm{radius}(E\cap{F}^{\bot}),$
@@
    where $\Delta_{J}^{s}$ stands for the $s$-dimensional regular simplex circumscribed around $\mathbf{B}^{s}.$
@@
    $\operatorname{inrad}\left(K\cap{F}^{\bot}\right)\leq\sqrt{\frac{d}{d-s}}\cdot\mathrm{radius}(E\cap{F}^{\bot}),$
@@
    where $\Box^{s}$ stands for the standard cube $[-1,1]^{s}.$
@@
This result shows that the section $K\cap F$ exhibits the same properties as a section through the center of the John ellipsoid. There is hardly any difference between the proofs of (2), (3) (and (S2), (S3)) and the corresponding statements for John ellipsoids. Everything reduces to the algebraic conditions on the projections of the contact vectors onto $F$. The proof of (1) (and (S1)) however, is more intricate.
We prove Theorems 1.4 and 1.6 in Sections 2 and 3. Then, in Section 4, we consider the problem of finding a homothety ratio such that enlarging the inner body in some sort of John position by this ratio, it will contain the outer body. In particular, we prove (3) in Theorem 1.7, and also show that no similar containment holds in the whole space $\mathbb{R}^{d}$. In Section 3.4, we state a certain converse of Theorems 1.4 and 1.6, that is, that the equations in this theorem ensure that $K$ is of locally maximum volume among certain affine images of $K$ under some additional conditions.
In Section 5, we discuss the problem of finding largest volume ellipsoids of revolution in a convex body, and in particular, we prove Theorem 1.7. We discuss the uniqueness of the solution to these John type problems in Section 6, and finally, in Section 7 outline how some of the results change for the dual question (often referred to as Löwner’s ellipsoid), where the smallest volume ellipsoid containing a convex body is considered. In Appendix A, we show that (1) of Theorem 1.7 is tight. Appendix B contains the proof of a technical fact that we use in Section 5.
@@ 1.4. Notation
We denote by $\mathbf{0}$ the origin of a vector space, and sometimes in a subscript, we specify the space, eg. $\mathbf{0}_{\mathcal{L}}$ in a space $\mathcal{L}$ of matrices. We use $\left\langle u,v\right\rangle$ to denote the standard inner product of the vectors $u,v\in\mathbb{R}^{d}$, and $\mathbf{B}^{d}$ for the closed Euclidean unit ball in $\mathbb{R}^{d}$ centered at the origin. We denote the positive hull of a set $S$ in $\mathbb{R}^{n}$ by $\operatorname{pos}\left(S\right)=\left\{\lambda x:\;\lambda\geq 0,x\in S\right\}.$ The polar of $S$ is defined as $S^{\circ}=\left\{p\in\mathbb{R}^{n}:\;\left\langle x,p\right\rangle\leq 1\,\text{ for all }x\in S\right\}$. Note that if $C$ is a positive cone (that is, the positive hull of some set), then $C^{\circ}=\left\{p\in\mathbb{R}^{n}:\;\left\langle x,p\right\rangle\leq 0\,\text{ for all }x\in C\right\}$. The group of orthogonal transformations on $\mathbb{R}^{d}$ is $O(d)$.
@@ 2.1. The normal and the tangent cone
$N_{C}\!\left(a\right)=\left\{p:\;\left\langle p,z-a\right\rangle\leq 0\quad\forall z\in C\right\}.$
@@
The following facts are well-known, and follow from standard separation properties of convex sets:
@@
Consider a convex set $C$ in $\mathbb{R}^{d}$ and a point $a\in\operatorname{cl}\left(C\right)$. Then
@@ 2.2. Operator spaces
We use $\mathcal{M}^{d}$ and $\mathcal{M}_{\mathrm{Sym}}^{d}$ to denote the space of operators and the space of self-adjoint operators from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, respectively. Define
@@
$\left\langle(M_{1},S_{1},g_{1}),(M_{2},S_{2},g_{2})\right\rangle_{\mathcal{R}}=\operatorname{tr}\left(M_{1}{M_{2}}^{\top}\right)+\operatorname{tr}\left(S_{1}{S_{2}}^{\top}\right)+\left\langle g_{1},g_{2}\right\rangle.$
@@ 2.3. A curve in $\mathcal{R}$
we have that $W$ is a skew-symmetric matrix, that is, $W=-{W}^{\top}$, and conversely, for every skew symmetric matrix $W$, there is a path $R_{t}$ in the orthogonal group satisfying (2.2).
@@
The first question is easy. We differentiate the determinant of the linear operator part of $A_{t}$.
@@
    $\left\langle(H,W,z),(v\otimes p,v\otimes p,p)\right\rangle_{\mathcal{R}}<0;$
@@
    $\left\langle(H,W,z),(v\otimes p,v\otimes p,p)\right\rangle_{\mathcal{R}}\leq 0;$
@@
$\left\langle\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}v,p\right\rangle=\left\langle\left(H+W\right)v+z,p\right\rangle=\left\langle Hv,p\right\rangle+\left\langle Wv,p\right\rangle+\left\langle z,p\right\rangle=$
$\left\langle(H,W,z),(v\otimes p,v\otimes p,p)\right\rangle_{\mathcal{R}}.$
@@
(2.6) $\left\langle(H,W,z),(v\otimes p,v\otimes p,p)\right\rangle_{\mathcal{R}}<0$
@@
$\mathrm{Cont}_{\mathcal{R}}=\left\{\left(v\otimes p,v\otimes p,p\right):\;v\in\mathrm{Cont}_{\mathbb{R}^{d}},p\in{L}^{\circ},\left\langle v,p\right\rangle=1\right\},$
@@
(2.7) $(H,W,z)\in\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right)$
@@
We fix a non-trivial subspace $F$ of $\mathbb{R}^{d}.$ We use $P_{F}$ to denote the orthogonal projection onto $F$. Set
$\mathcal{Q}_{F}=\left\{Q\in\operatorname{Lin}\left(\mathbb{R}^{d}\right):\;P_{F}QP_{F}=\mathbf{0}_{F},\operatorname{rank}\left(P_{{F}^{\bot}}QP_{{F}^{\bot}}\right)=d-\dim{F},\operatorname{tr}\left(P_{{F}^{\bot}}QP_{{F}^{\bot}}\right)=0\right\},$
@@
where $\mathbf{0}_{s\times s}$ is the $s\times s$ zero matrix, $Q_{1}$ and ${Q_{2}}^{\top}$ are $s\times(d-s)$ matrices, and $Q_{3}$ is a square matrix of rank $d-s$ of zero trace.
@@
3.1. Slight variants of Theorem 1.6 and 1.4
In this subsection, we prove variants of these two theorems without a translation $z$, and as a price, we do not obtain (1.3).
The following result differs slightly from Theorem 1.6.
Theorem 3.2.
Let $K$, $L$ and $F$ satisfy the conditions of Theorem 1.6, and assume that the origin is in the interior of $K$, and that $K$ is of locally maximum volume among elements of $\mathcal{R}_{F}\!\left(K\right)$ contained in $L$.
Then we have contact points of $K$ and $L$ and weights satisfying (1.1), (1.2), (1.4), and (1.5).
@@
Thus, there are matrices $Q\in\mathcal{Q}_{F}$ and $S\in\mathcal{M}_{\mathrm{Sym}}^{d}$ such that $\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q,S,\mathbf{0}\right)\in\mathcal{C}_{\mathcal{R}}$. That is, there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}>0$ such that $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and the matrix $M=\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}$ satisfies the conditions
@@
The following result differs slightly from Theorem 1.4.
Theorem 3.3.
Let $K$, $L$ and $F$ satisfy the conditions of Theorem 1.4, and assume that the origin is in the interior of $L$, and that $K$ is of locally maximum volume among elements of $\mathcal{A}_{F}\!\left(K\right)$ contained in $L$.
Then we have contact points of $K$ and $L$ and weights satisfying (1.1), (1.2), and (1.4).
The proof is essentially identical to the previous one.
@@
Thus, there are matrices $Q\in\mathcal{Q}_{F}$ and $S\in\mathcal{M}^{d}$ such that $\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q,S,\mathbf{0}\right)\in\mathcal{C}_{\mathcal{R}}$. That is, there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}>0$ such that $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and the matrix $M=\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}$ satisfies the conditions

### v2

@@ Abstract.
Finding a largest Euclidean ball in a given convex body $K\subset\mathbb{R}^{d}$ and finding a largest volume ellipsoid in $K$ are two problems of fundamentally different nature. The first is a purely Euclidean problem, where we consider scaled copies of the origin-centered closed unit ball, whereas in the second problem, we search among all affine copies of the unit ball.
In this paper, we interpolate between these two classical problems by considering ellipsoids of revolution. More generally, we study pairs of convex bodies $K$ and $L$, and seek a largest-volume affine image of $K$ contained within $L$, subject to certain restrictions on the allowed affine transformations. We derive first-order necessary conditions for optimality, generalizing known conditions from the unrestricted affine setting.
Using these conditions, we show that an extremal ellipsoid of revolution exhibits properties analogous to those of either the largest-volume ellipsoid or the largest Euclidean ball, depending on whether the ellipsoid is considered along its axis of revolution or along the orthogonal complement of that axis.
@@ 1.2. A more general problem
A natural generalization of this question was studied by V. Milman [TJ89, Theorem 14.5], Lewis [Lew79], Giannopoulos, Perissinaki and Tsolomitis [GPT01], Bastero and Romance [BR02], and Gordon, Litvak, Meyer and Pajor [GLMP04]: given convex bodies $K$ and $L$ in $\mathbb{R}^{d}$, find a maximal volume affine image of $K$ that is contained in $L$. In this work, we study the same problem under additional restrictions on the affine transformation.
@@
Note a subtle but very important difference between this general case and ellipsoids. To define ellipsoids of revolution with an $s$-dimensional axis, we only need to specify the dimension $s$ of the axis, but not the axis $F$. In the general case introduced here, we fix $F$, apply an $F$-operator $A$ and then allow a rotation $R$. The advantage of this is that in this way, we can specify what ‘part’ of $K$ we allow to be changed by an affine transformation (the $F$-component), and what part (the ${F}^{\bot}$ component) is rigid, that is, it only allows similarities. A simple illustration is the task of finding the largest area isosceles triangle contained in a planar set $L$. In our framework, we simply take $K$ to be a regular triangle, and $F$ to be one of its altitude lines. Then $\mathcal{R}_{F}\!\left(K\right)$ is exactly the set of isosceles triangles.
When $F=\{\mathbf{0}\}$, then Problem 4 is the problem discussed at the beginning of this subsection considered by Lewis, Milman and others. On the other hand, when $F=\mathbb{R}^{d}$, then we are looking for the largest similar copy of $K$ in $L$.
@@ 1.3. Results
[…]
@@
Let $K,L\subset\mathbb{R}^{d}$ be convex bodies containing the origin in their interiors and such that $K\subseteq L$, and let $F$ be a linear subspace in $\mathbb{R}^{d}$. Assume that $K$ is a solution to John’s general problem with fixed axis $F$ for $K$ and $L$, that is, Problem 3. Then there are
@@
    contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L$ and $K$,
@@
(1.1) $P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}\right)P_{F}=P_{F};$
@@
(1.3) $\operatorname{tr}\left(\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}\right)=\sum\limits_{i\in[m]}\upalpha_{i}=d.$
As a corollary, we will obtain the following.
Theorem 1.5 (Containment for John’s general problem with fixed axis).
In addition to the assumptions of Theorem 1.4, suppose that $P_{F}K\subset L\cap F$. Then there exists a translation vector $z$ in the relative interior of $P_{F}K$ such that
$P_{F}(K-z)\subset\left(L-z\right)\cap F\subset-d\cdot P_{F}(K-z).$
In particular, such a translation vector $z$ exists whenever $K$ is symmetric with respect to $F$.
In the case $F=\mathbb{R}^{d}$, Theorem 1.4 and Theorem 1.5 were obtained in [GLMP04].
@@
With the notation of Theorem 1.4, assume that $K$ contains the origin in its interior and is a solution to John’s general problem with rotated axis for $K$, $L$ and the subspace $F$. Then there are contact pairs of $L$ and $K$, and weights as in Theorem 1.4 such that in addition to identities (1.1) – (1.3), we also have
(1.4) $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}=\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}.$
@@
Let an origin-centered ellipsoid $E$ be a solution to Problem 1 for a convex body $K$ and a linear subspace $F$ in $\mathbb{R}^{d}$ of dimension $s\in\{0,1,\dots,d-1\}$. Then
@@
    $\operatorname{inrad}\left(K\cap{F}^{\bot}\right)\leq\frac{d}{d-s}\cdot\mathrm{radius}(E\cap{F}^{\bot});$
@@
    where $\Delta_{J}^{s}$ stands for the $s$-dimensional regular simplex circumscribed around $\mathbf{B}^{s}$;
@@
    $\operatorname{inrad}\left(K\cap{F}^{\bot}\right)\leq\sqrt{\frac{d}{d-s}}\cdot\mathrm{radius}(E\cap{F}^{\bot});$
@@
    where $\Box^{s}$ stands for the standard cube $[-1,1]^{s}$;
@@
Note that (3) and (S3) are sharp: $\mathbf{B}^{d}$ is the John ellipsoid of $\Delta^{d}$ (resp., of $\Box^{d}$), and hence, it is the solution of John’s problem for ellipsoids with a fixed axis of revolution for any $F$. If we now choose $F$ properly, then we have equality in these two inequalities.
On the other hand, (2) and (S2) are identical to the best bounds that are known for sections of $K$ and its (unrestricted) John ellipsoid. See our discussion of optimality in Subsection 5.5.
There is hardly any difference between the proofs of (2), (3) (and (S2), (S3)) and the corresponding statements for John ellipsoids. Everything reduces to the algebraic conditions on the projections of the contact vectors onto $F$. The proof of (1) (and (S1)) however, is more intricate.
We prove Theorems 1.4 and 1.6 in Sections 2 and 3. In Subsection 3.3, we state a certain converse of Theorems 1.4 and 1.6, that is, that the equations in this theorem ensure that $K$ is of locally maximum volume among certain affine images of $K$ under some additional conditions.
In Section 4, we consider the problem of finding a homothety ratio such that enlarging the inner body in some sort of John position by this ratio, it will contain the outer body. In particular, we prove (3) in Theorem 1.7, and also show that no similar containment holds in the whole space $\mathbb{R}^{d}$.
We may refer to the notion of an isosceles triangle as a ‘semi-Euclidean’ notion, since it is neither an affine notion nor purely Euclidean – that is, a notion invariant under similarity alone. In Subsection 4.4, we show an application of our results to isosceles triangles, and more generally, to right cones to illustrate how our language is capable of addressing semi-Euclidean questions.
In Section 5, we discuss the problem of finding largest volume ellipsoids of revolution in a convex body, and in particular, we prove Theorem 1.7. We discuss the uniqueness of the solution to these John type problems in Section 6, and finally, in Section 7, we outline how some of the results change for the dual question (often referred to as Löwner’s ellipsoid), where the smallest volume ellipsoid containing a convex body is considered. In Appendix A, we show that (1) of Theorem 1.7 is tight. Appendix B contains the proof of a technical fact that we use in Section 5.
@@ 1.4. Notation
We denote by $\mathbf{0}$ the origin of a vector space, and sometimes in a subscript, we specify the space, e.g. $\mathbf{0}_{\mathcal{L}}$ in a space $\mathcal{L}$ of matrices. We use $\left\langle u,v\right\rangle$ to denote the standard inner product of the vectors $u,v\in\mathbb{R}^{d}$, and $\mathbf{B}^{d}$ for the closed Euclidean unit ball in $\mathbb{R}^{d}$ centered at the origin. We denote the positive hull of a set $S$ in $\mathbb{R}^{n}$ by $\operatorname{pos}\left(S\right)=\left\{\lambda x:\;\lambda\geq 0,x\in S\right\}.$ The polar of a set $S$ is defined as $S^{\circ}=\left\{p\in\mathbb{R}^{n}:\;\left\langle x,p\right\rangle\leq 1\,\text{ for all }x\in S\right\}$. Note that for a positive cone $C$ (that is, the positive hull of some set), we have $C^{\circ}=\left\{p\in\mathbb{R}^{n}:\;\left\langle x,p\right\rangle\leq 0\,\text{ for all }x\in C\right\}$. The group of orthogonal transformations on $\mathbb{R}^{d}$ is $O(d)$.
@@ 2.1. The normal and the tangent cone
$N_{C}\!\left(a\right)=\left\{p\in\mathbb{R}^{n}:\;\left\langle p,z-a\right\rangle\leq 0\quad\forall z\in C\right\}.$
@@
The following facts are well-known, and follow from standard separation properties of convex sets [HUL04, Section 5.2]:
@@
Consider a convex set $C$ in $\mathbb{R}^{n}$ and a point $a\in\operatorname{cl}\left(C\right)$. Then
@@ 2.2. Operator spaces
We use $\mathcal{M}^{d}$ and $\mathcal{M}_{\mathrm{Sym}}^{d}$ to denote the space of operators and the space of self-adjoint operators from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$, respectively. The inner product on $\mathcal{M}^{d}$ is defined as $\left\langle A,B\right\rangle_{\mathcal{M}^{d}}=\operatorname{tr}\left(A{B}^{\top}\right)$.
Define
@@
$\left\langle(M_{1},S_{1},g_{1}),(M_{2},S_{2},g_{2})\right\rangle_{\mathcal{R}}=\left\langle M_{1},M_{2}\right\rangle_{\mathcal{M}^{d}}+\left\langle S_{1},S_{2}\right\rangle_{\mathcal{M}^{d}}+\left\langle g_{1},g_{2}\right\rangle.$
@@ 2.3. A curve in $\mathcal{R}$
we have that $W$ is a skew-symmetric matrix, that is, $W=-{W}^{\top}$, and conversely, for every skew-symmetric matrix $W$, there is a path $R_{t}$ in the orthogonal group satisfying (2.2).
@@
The first question is easy. We differentiate the determinant of the linear operator part of $A_{t}$:
@@
    $\left\langle(H,W,z),(p\otimes v,p\otimes v,p)\right\rangle_{\mathcal{R}}<0;$
@@
    $\left\langle(H,W,z),(p\otimes v,p\otimes v,p)\right\rangle_{\mathcal{R}}\leq 0;$
@@
$\left\langle\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}v_{t},p\right\rangle=\left\langle\left(H+W\right)v+z,p\right\rangle=\left\langle Hv,p\right\rangle+\left\langle Wv,p\right\rangle+\left\langle z,p\right\rangle=$
$\left\langle(H,W,z),(p\otimes v,p\otimes v,p)\right\rangle_{\mathcal{R}}.$
@@
(2.6) $\left\langle(H,W,z),(p\otimes v,p\otimes v,p)\right\rangle_{\mathcal{R}}<0$
@@
$\mathrm{Cont}_{\mathcal{R}}=\left\{\left(p\otimes v,p\otimes v,p\right):\;v\in\mathrm{Cont}_{\mathbb{R}^{d}},p\in{L}^{\circ},\left\langle v,p\right\rangle=1\right\},$
@@
(2.7) $(H,W,z)\in\mathrm{int}\left({\mathcal{C}_{\mathcal{R}}}^{\circ}\right),$
@@
We fix a non-trivial subspace $F$ of $\mathbb{R}^{d}$ of dimension $s.$ We use $P_{F}$ to denote the orthogonal projection onto $F$. Set
$\mathcal{Q}_{F}=\left\{Q\in\operatorname{Lin}\left(\mathbb{R}^{d}\right):\;P_{F}QP_{F}=\mathbf{0}_{F},\operatorname{rank}\left(P_{{F}^{\bot}}QP_{{F}^{\bot}}\right)=d-s,\operatorname{tr}\left(P_{{F}^{\bot}}QP_{{F}^{\bot}}\right)=0\right\},$
@@
where $\mathbf{0}_{s\times s}$ is the $s\times s$ zero matrix, $Q_{1}$ and ${Q_{2}}^{\top}$ are $s\times(d-s)$ matrices, and $Q_{3}$ is a full rank $(d-s)\times(d-s)$ matrix of zero trace.
@@
3.1. Proofs of Theorem 1.6 and 1.4
@@
Thus, there are matrices $Q\in\mathcal{Q}_{F}$ and $S\in\mathcal{M}_{\mathrm{Sym}}^{d}$ such that $\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q,S,\mathbf{0}\right)\in\mathcal{C}_{\mathcal{R}}$. That is, there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}>0$ such that $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and the matrix $M=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}$ satisfies the conditions
@@
The proof of Theorem 1.4 is essentially identical to the previous one.
@@
Thus, there are matrices $Q\in\mathcal{Q}_{F}$ and $S\in\mathcal{M}^{d}$ such that $\left(\mathrm{Id}_{\mathbb{R}^{d}}+Q,S,\mathbf{0}\right)\in\mathcal{C}_{\mathcal{R}}$. That is, there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}>0$ such that $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and the matrix $M=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}$ satisfies the conditions